# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ Abstract
We propose SDORE, a semi-supervised deep Sobolev regressor, for the nonparametric estimation of the underlying regression function and its gradient. SDORE employs deep neural networks to minimize empirical risk with gradient norm regularization, allowing computation of the gradient norm on unlabeled data. We conduct a comprehensive analysis of the convergence rates of SDORE and establish a minimax optimal rate for the regression function. Crucially, we also derive a convergence rate for the associated plug-in gradient estimator, even in the presence of significant domain shift. These theoretical findings offer valuable prior guidance for selecting regularization parameters and determining the size of the neural network, while showcasing the provable advantage of leveraging unlabeled data in semi-supervised learning. To the best of our knowledge, SDORE is the first provable neural network-based approach that simultaneously estimates the regression function and its gradient, with diverse applications including nonparametric variable selection and inverse problems. The effectiveness of SDORE is validated through an extensive range of numerical simulations and real data analysis.
62G05,
62G08,
65N21,
Nonparametric regression,
gradient estimation,
variable selection,
convergence rate,
gradient penalty,
deep neural network,
keywords:
[class=MSC]
keywords:
[…]
, , and
1 Introduction
Nonparametric regression plays a pivotal role in both statistics and machine learning, possessing an illustrious research history as well as a vast compendium of related literature (Györfi et al., 2002; Wasserman, 2006; Tsybakov, 2009). Let $\Omega\subseteq{\mathbb{R}}^{d}$, $d\geq 1$, be a bounded and connected domain with sufficiently smooth boundary $\partial\Omega$. Consider the following nonparametric regression model
@@ 1 Introduction
where $Y\in{\mathbb{R}}$ is the response associated with the covariate $X\in\Omega$ and $f_{0}$ is the unknown regression function. Here $\xi$ represents a random noise term satisfying ${\mathbb{E}}[\xi|X]=0$ and ${\mathbb{E}}[\xi^{2}|X]<\infty$. The primary task of nonparametric regression involves estimating the conditional expectation $f_{0}(x)$ of the response $Y$, given a covariate vector $X=x$, an endeavor typically achieved through empirical least-squares risk minimization:
@@
where $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are independently and identically distributed random copies of $(X,Y)$, and ${\mathcal{F}}$ is a pre-specific hypothesis class, such as deep neural network class in this paper. However, the conventional empirical risk minimizer often falls short in fulfilling certain desired requirements. For instance, constructing estimators resilient to input perturbations requires constraints on the norm of estimator gradient.
To address these challenges, Sobolev regularization (gradient norm penalty) was introduced in deep learning by Drucker and Le Cun (1991); Drucker and Le Cun (1992):
@@
where $\lambda>0$ is the regularization parameter. Substantial numeral experiments have consistently demonstrated that the imposition of a gradient penalty contributes to the enhancement of the stability and generalization of deep learning models. The strategy surrounding gradient penalty was adopted by Rifai et al. (2011) as a technique to learn robust features using auto-encoders. This method was further utilized to augment the stability of deep generative models as highlighted in the work of Nagarajan and Kolter (2017); Roth et al. (2017); Gulrajani et al. (2017); Gao et al. (2022). Significantly, the gradient norm, being a local measure of sensitivity to input perturbations, has seen a plethora of research focusing on its use for adversarial robust learning. This is reflected in studies conducted by Lyu, Huang and Liang (2015); Hein and Andriushchenko (2017); Ororbia II, Kifer and Giles (2017); Novak et al. (2018); Jakubovitz and Giryes (2018); Ross and Doshi-Velez (2018).
@@
Simultaneously estimation the regression function and its of gradient (derivatives) carries a wide span of applications across various fields, including the factor demand and cost estimation in economics (Shephard, 1981), trend analysis for time series data (Rondonotti, Marron and Park, 2007), the analysis of human growth data (Ramsay and Silverman, 2002), and the modeling of spatial process (Banerjee, Gelfand and Sirmans, 2003). Furthermore, estimating gradient plays a pivotal role in the modeling of functional data (Müller and Yao, 2010; Dai, Müller and Tao, 2018), variable selection in nonparametric regression Rosasco et al. (2010); Mosci et al. (2012); Rosasco et al. (2013), and inverse problems (Hu, Shu and Zou, 2018). There are four classical approaches to nonparametric gradient estimation: local polynomial regression (De Brabanter et al., 2013), smoothing splines (Heckman and Ramsay, 2000), kernel ridge regression (Liu and Li, 2023), and difference quotients (Müller, Stadtmüller and Schmitt, 1987). However, local polynomial regression and smoothing spline regression are only applicable to fixed-design setting and low-dimensional problems. The generalization of these methodologies to address high-dimensional problems is met with a significant challenge popularly known as the computational curse of dimensionality (Bellman, 1961; Wasserman, 2006; Goodfellow, Bengio and Courville, 2016). This phenomenon refers to the fact that the computational complexity can increase exponentially with dimension. In contrast, deep neural network-based methods, which are mesh-free, exhibit direct applicability to high-dimensional problems, providing a solution to mitigate this inherent challenge. The plug-in kernel ridge regression estimators have demonstrated applicability for estimating derivatives across both univariate and multivariate regressions within a random-design setting (Liu and Li, 2023; Liu, Li and Li, 2023). However, these estimators present certain inherent limitations compared to deep neural networks. From a computational complexity standpoint, the scale of the kernel grows quadratically or even cubically with the number of samples. In contrast, deep neural networks exhibit the ability to handle larger datasets, especially when deployed on modern hardware architectures.
@@
Recently, there has been a substantial literature outlining the convergence rates of deep nonparametric regression (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell, Liang and Misra, 2021; Kohler, Krzyżak and Langer, 2022; Jiao et al., 2023). However, the theoretical foundation of Sobolev regularized least-squares using deep neural networks remains relatively underdeveloped. Consequently, two fundamental questions need to be addressed:
@@
plays a pivotal role in nonparametric regression and has been investigated by Györfi et al. (2002). We establish non-asymptotic convergence rates for the deep Sobolev regressor and demonstrate that the norm of its gradient is uniformly bounded, shedding light on the considerable stability and favorable generalization properties of the estimator. Furthermore, under certain mild conditions, we derive non-asymptotic convergence rates for the plug-in derivative estimator based on SDORE. This illustrates how abundant unlabeled data used in SDORE (3) improves the performance of the standard gradient penalized regressor (2). We subsequently apply SDORE to nonparametric variable selection and inverse problems. The efficacy of this method is substantiated through numerous numerical examples.
1.1 Contributions
@@
    We introduce a novel semi-supervised deep estimator within the framework of Sobolev penalized regression. A large amount of unlabeled data is employed to estimate the Sobolev penalty term. We demonstrate that this deep neural network-based estimator achieves the minimax optimal rate (Theorem 4.3). Meanwhile, with the appropriate selection of the regularization parameter, the norm of the estimator’s gradient can be uniformly bounded, thereby illustrating its remarkable stability and generalization capacities from a theoretical standpoint.
@@
    Under certain mild conditions, we establish an oracle inequality for gradient estimation using the plug-in deep Sobolev regressor (Lemma 5.2). Notably, this oracle inequality is applicable to any convex hypothesis class. This represents a significant theoretical advancement beyond existing nonparametric plug-in gradient estimators, which are based on linear approximation (Stone, 1985; Liu and Li, 2023), by extending the framework to handle more complex hypothesis classes involved in nonlinear approximation (DeVore and Lorentz, 1993). Furthermore, we derive a convergence rate for the gradient of the deep neural network-based estimator, providing valuable a priori guidance for selecting regularization parameters and choosing the size of the neural network (Theorem 5.4).
@@
    We derive a convergence rate for semi-supervised estimator (Theorem 5.6), which sheds light on the quantifiable advantages of incorporating unlabeled data into the supervised learning. This improvement is actualized under the condition that density ratio between the marginal distribution of the labeled data and the distribution of the unlabeled data remains uniformly bounded. This novel finding promises to enrich our theoretical comprehension of semi-supervised learning, particularly in the context of deep neural networks.
@@
    The gradient estimator introduces a novel tool with potential applications in areas such as nonparametric variable selection and inverse problems in partial differential equations. In the case where the regression function exhibits sparsity structure (Assumption 7), we prove that the convergence rate depends only on the number of relevant variables, rather than the data dimension (Corollary 6.2). Moreover, we establish the selection consistency of the deep Sobolev regressor (Corollary 6.3), showing that, with a sufficiently large number of labeled data pairs, the estimated relevant set is highly likely to match the ground truth relevant set. In addition, we derive the convergence rate for inverse source problems with interior measurements (Corollary 6.4). To validate our approach, we conduct a series of numerical experiments, which confirm the effectiveness and reliability of our proposed methodology.
1.2 Main Results Overview
@@
The main theoretical results derived in this paper are summarized in Table 1. We utilize the deep Sobolev regressor to tackle two distinct application scenarios: nonparametric variable selection and inverse source problems involving partial differential equations. We present the theoretical findings related to nonparametric variable selection in Table 2, including the convergence rate and selection consistency. Additionally, for the inverse source problems,the $H^{1}$-norm rate for the recovered solution $\hat{u}_{{\mathcal{D}}}^{\lambda}$ and the $H^{1}$-dual-norm rate for the recovered source term $\hat{f}_{{\mathcal{D}}}^{\lambda}$ are also presented in Table 2.
Table 1: Convergence Rates for Sobolev Penalized Estimators
Estimator Reg. Param. Convergence Rates
DORE [^1] Deep Sobolev regressor defined in (8). $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}})$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+2s}})$ Theorem 4.3
${\mathbb{E}}\|\nabla\hat{f}_{{\mathcal{D}}}^{\lambda}\|^{2}$ ${\mathcal{O}}(1)$
DORE [^2] Deep Sobolev regressor defined in (11). $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}})$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+4s}})$ Theorem 5.4
${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d+4s}})$
SDORE [^3] Semi-supervised deep Sobolev regressor defined in (11). $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}})$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+4s}})+{\mathcal{O}}(n^{\frac{d}{2(d+4s)}}m^{-\frac{1}{2}})$ Theorem 5.6
${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d+4s}})+{\mathcal{O}}(n^{\frac{d+2s}{2(d+4s)}}m^{-\frac{1}{2}})$
Table 2: Theoretical Results for Applications
@@
$\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}})$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d^{*}+4s}})$ Corollary 6.2
${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}})$
Selection consistency Rel. Set ${\mathcal{I}}(f_{0})$ $\lim_{n\rightarrow\infty}\operatorname{Pr}\{{\mathcal{I}}(f_{0})={\mathcal{I}}(\hat{f}_{{\mathcal{D}}}^{\lambda})\}=1$ Corollary 6.3
@@
Inverse Source Problem
Reg. Param. Recon. Error Convergence Rates
$\lambda={\mathcal{O}}((1+\sigma^{\frac{3s}{d+4s}})n^{-\frac{s}{d+4s}})$ ${\mathbb{E}}\|\hat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|$ ${\mathcal{O}}((1+\sigma^{\frac{3s}{d+4s}})n^{-\frac{s}{d+4s}})$ Corollary 6.4
${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{*}\|$ ${\mathcal{O}}((1+\sigma^{\frac{3s}{2(d+4s)}})n^{-\frac{s}{2(d+4s)}})$
1.3 Preliminaries

### v2

# Semi-Supervised Deep Sobolev Regression: Estimation and Variable Selection by ReQU Neural Network

@@ Abstract
We propose SDORE, a Semi-supervised Deep SObolev REgressor, for the nonparametric estimation of the underlying regression function and its gradient. SDORE employs deep ReQU neural networks to minimize the empirical risk with gradient norm regularization, allowing the approximation of the regularization term by unlabeled data. Our study includes a thorough analysis of the convergence rates of SDORE in $L^{2}$-norm, achieving the minimax optimality. Further, we establish a convergence rate for the associated plug-in gradient estimator, even in the presence of significant domain shift. These theoretical findings offer valuable insights for selecting regularization parameters and determining the size of the neural network, while showcasing the provable advantage of leveraging unlabeled data in semi-supervised learning. To the best of our knowledge, SDORE is the first provable neural network-based approach that simultaneously estimates the regression function and its gradient, with diverse applications such as nonparametric variable selection. The effectiveness of SDORE is validated through an extensive range of numerical simulations.
Index Terms:
Nonparametric regression, gradient estimation, variable selection, convergence rate, gradient penalty, deep neural network
I Introduction
Nonparametric regression plays a pivotal role in both statistics and machine learning, possessing an illustrious research history as well as a vast compendium of related literature [1, 2, 3]. Let $\Omega\subseteq{\mathbb{R}}^{d}$, $d\geq 1$, be a bounded and connected domain with sufficiently smooth boundary $\partial\Omega$. Consider the following nonparametric regression model
@@ I Introduction
where $Y\in{\mathbb{R}}$ is the response associated with the covariate $X\in\Omega$, and $f_{0}$ is the unknown regression function. Here $\xi$ represents a random noise term satisfying ${\mathbb{E}}[\xi|X]=0$ and ${\mathbb{E}}[\xi^{2}|X]<\infty$. The primary task of nonparametric regression involves estimating the conditional expectation $f_{0}(x)$ of the response $Y$, given a covariate $X=x$. This estimation is typically achieved through empirical least-squares risk minimization:
@@
where $\{(X_{i},Y_{i})\}_{i=1}^{n}$ is a set of independently and identically distributed random copies of $(X,Y)$, and ${\mathcal{F}}$ is a pre-specific hypothesis class, such as deep ReQU neural network class in this paper. While empirical least-squares risk minimization is straightforward to implement and comes with solid theoretical guarantees, it does not fully meet all desired criteria. One major drawback is that the method places no constraints on the gradient of the estimator, allowing for the possibility of an arbitrarily large gradient norm. This can make the least-squares estimator highly sensitive to the input perturbations. Furthermore, while the least-squares estimator ensures convergence in terms of function values, the convergence in terms of derivatives can not be guaranteed.
To address these challenges, Sobolev regularization, also known as gradient penalty, was introduced in deep learning by [4, 5]:
@@
where $\lambda>0$ is the regularization parameter, and $D_{k}f$ denotes the partial derivative of $f$ with respect to the $k$-th input variable. Substantial numerical experiments have consistently demonstrated that the imposition of a gradient penalty contributes to the enhancement of the stability and generalization of deep learning models. The strategy surrounding gradient penalty was adopted by [6] as a technique to learn robust features using auto-encoders. This method was further utilized to augment the stability of deep generative models as highlighted in the work of [7, 8, 9, 10]. Significantly, the gradient norm, being a local measure of sensitivity to input perturbations, has seen a plethora of research focusing on its use for adversarial robust learning. This is reflected in studies conducted by [11, 12, 13, 14, 15, 16].
@@
Simultaneously estimation the regression function and its of gradient (derivatives) carries a wide span of applications across various fields, including the factor demand and cost estimation in economics [17], trend analysis for time series data [18], the analysis of human growth data [19], and the modeling of spatial process [20]. Furthermore, estimating gradient plays a pivotal role in the modeling of functional data [21, 22], variable selection in nonparametric regression [23, 24, 25], and inverse problems [26]. There are four classical approaches to nonparametric gradient estimation: local polynomial regression [27], smoothing splines [28], kernel ridge regression [29], and difference quotients [30]. However, local polynomial regression and smoothing spline regression are only applicable to fixed-design setting and low-dimensional problems. The generalization of these methodologies to address high-dimensional problems is met with a significant challenge popularly known as the computational curse of dimensionality [31, 2, 32]. This phenomenon refers to the fact that the computational complexity can increase exponentially with dimension. In contrast, deep neural network-based methods, which are mesh-free, exhibit direct applicability to high-dimensional problems, providing a solution to mitigate this inherent challenge. The plug-in kernel ridge regression estimators have demonstrated applicability for estimating derivatives across both univariate and multivariate regressions within a random-design setting [29, 33]. However, these estimators present certain inherent limitations compared to deep neural networks. From a computational complexity standpoint, the scale of the kernel grows quadratically or even cubically with the number of samples. In contrast, deep neural networks exhibit the ability to handle larger datasets, especially when deployed on modern hardware architectures.
@@
Recently, there has been a substantial literature outlining the convergence rates of deep nonparametric regression [34, 35, 36, 37, 38, 39, 40]. However, the theoretical foundation of Sobolev regularized least-squares using deep neural networks remains relatively underdeveloped. Consequently, two fundamental questions need to be addressed:
@@
plays a pivotal role in nonparametric regression and has been investigated by [41, 42, 43, 1]. We establish non-asymptotic convergence rates for the deep Sobolev regressor and demonstrate that the norm of its gradient is uniformly bounded, shedding light on the considerable stability and favorable generalization properties of the estimator. Furthermore, under certain mild conditions, we derive non-asymptotic convergence rates for the plug-in derivative estimator based on SDORE. This illustrates how abundant unlabeled data used in SDORE (3) improves the performance of the standard gradient penalized regressor (2). We subsequently apply SDORE to nonparametric variable selection. The efficacy of this method is substantiated through numerous numerical examples.
I-A Contributions
@@
    We introduce a novel semi-supervised deep estimator within the framework of Sobolev penalized regression. A large amount of unlabeled data is employed to estimate the Sobolev penalty term. We demonstrate that this deep ReQU neural network-based estimator achieves the minimax optimal rate (Theorem IV.3). Meanwhile, with the appropriate selection of the regularization parameter, the norm of the estimator’s gradient can be uniformly bounded, thereby illustrating its remarkable stability and generalization capacities from a theoretical standpoint.
@@
    Under certain mild conditions, we establish an oracle inequality for gradient estimation using the plug-in deep Sobolev regressor (Lemma V.2). Notably, this oracle inequality is applicable to any convex hypothesis class. This represents a significant theoretical advancement beyond existing nonparametric plug-in gradient estimators, which are based on linear approximation [44, 29], by extending the framework to handle more complex hypothesis classes involved in nonlinear approximation [45]. Furthermore, we derive a convergence rate for the gradient of the deep ReQU neural network-based estimator, providing valuable a priori guidance for selecting regularization parameters and choosing the size of the neural network (Theorem V.4).
@@
    We derive a convergence rate for semi-supervised estimator (Theorem V.6), which sheds light on the quantifiable advantages of incorporating unlabeled data into the supervised learning. This improvement is actualized under the condition that density ratio between the marginal distribution of the labeled data and the distribution of the unlabeled data remains uniformly bounded. This novel finding promises to enrich our theoretical comprehension of semi-supervised learning, particularly in the context of deep neural networks.
@@
    The gradient estimator introduces a novel tool with potential applications in areas such as nonparametric variable selection. In the case where the regression function exhibits sparsity structure (Assumption 7), we prove that the convergence rate depends only on the number of relevant variables, rather than the data dimension (Corollary VI.1). Moreover, we establish the selection consistency of the deep Sobolev regressor (Corollary VI.2), showing that, with a sufficiently large number of labeled data pairs, the estimated relevant set is highly likely to match the ground truth relevant set. To validate our approach, we conduct a series of numerical experiments, which confirm the effectiveness and reliability of our proposed methodology.
I-B Main Results Overview
In this work, we focus on two estimators in the setting of nonparametric regression (1). The Deep SObolev REgressor (DORE) is derived from the regularized empirical risk minimization:
$\widehat{f}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}\\ +\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2},$ (4)
where ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is a set of independent copies of $(X,Y)$, $\lambda>0$ is the regularization parameter, and ${\mathcal{F}}$ is a class of deep ReQU neural networks. In some application scenarios, the regularization term in (4) is intractable analytically. To address this issue, we approximate the regularization term by its data-driven counterpart, yielding the following semi-supervised empirical risk minimizer
$\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}\\ +\frac{\lambda}{m}\sum_{i=1}^{m}\sum_{k=1}^{d}|D_{k}f(Z_{i})|^{2},$ (5)
where ${\mathcal{S}}=\{Z_{i}\}_{i=1}^{m}$ is a set of independently and identically random variables drawn from $\nu_{X}$.
The main theoretical results derived in this paper are summarized in Table I. As shown in Theorem IV.3, the convergence rate of the deep Sobolev regressor in $L^{2}$-norm achieves the minimax optimality. However, Theorems V.6 and V.4 demonstrate that the convergence rates in $L^{2}$-norm and $H^{1}$-semi-norm is sub-optimal.
TABLE I: Convergence Rates for Sobolev Penalized Estimators
Estimator Reg. Param. Convergence Rates Minimax Opt.
DORE $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ Theorem IV.3
Def. (11) ${\mathbb{E}}\|\nabla\hat{f}_{{\mathcal{D}}}^{\lambda}\|^{2}$ ${\mathcal{O}}(1)$
DORE $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+4s}}\log^{4}n)$ Theorem V.4
Def. (14) ${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$
SDORE $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d+4s}}\log^{4}n)+{\mathcal{O}}(n^{\frac{d}{d+4s}}\log^{4}nm^{-1})$ Theorem V.6
Def. (14) ${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)+{\mathcal{O}}(n^{\frac{d+s}{d+4s}}\log^{2}nm^{-1})$
@@ I-B Main Results Overview
We utilize the deep Sobolev regressor to tackle an application scenarios: nonparametric variable selection. We present the theoretical findings related to nonparametric variable selection in Table II, including the convergence rate and selection consistency.
TABLE II: Theoretical Results for Applications
@@
$\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}}\log^{2}n)$ ${\mathbb{E}}\|\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|^{2}$ ${\mathcal{O}}(n^{-\frac{2s}{d^{*}+4s}}\log^{4}n)$ Corollary VI.1
${\mathbb{E}}\|\nabla(\hat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|^{2}$ ${\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}}\log^{2}n)$
Selection consistency Rel. Set ${\mathcal{I}}(f_{0})$ $\lim_{n\rightarrow\infty}\operatorname{Pr}\{{\mathcal{I}}(f_{0})={\mathcal{I}}(\hat{f}_{{\mathcal{D}}}^{\lambda})\}=1$ Corollary VI.2
@@
I-C Preliminaries and notations